OpenStudy (anonymous):

more implicit differentials for what values of x does the curve y^2 -x^4 + 2xy -18x^2 = 10 have vertical tangent lines?

OpenStudy (anonymous):

@ganeshie8 @IrishBoy123 any ideas?

zepdrix (zepdrix):

Horizontal tangent lines when: \(\large\rm y'=0\) Vertical tangent lines when: \(\large\rm y'=\frac{stuff}{0}\) Have you tried finding your y' yet? :)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i feel i messed up somewhere though so im redoing that right now

OpenStudy (anonymous):

im getting \[y'= \frac{ 4x(x^2+9) }{ 2(y+x) }\]

zepdrix (zepdrix):

Hmm I think there is another term on top. Did you forget to product rule again? :)\[\large\rm y^2 -x^4 + \color{orangered}{2xy} -18x^2 = 10\]

OpenStudy (anonymous):

no i just simplified from 4x^3 + 36x

zepdrix (zepdrix):

But where is the -2y in the numerator? :o hmm

OpenStudy (anonymous):

oh your totally right

OpenStudy (anonymous):

argh your so clever :P

zepdrix (zepdrix):

\[\large\rm y^2 -x^4 + 2xy -18x^2 = 10\]Differentiating gives,\[\large\rm 2yy'-4x^3+2y+2xy'-36x=0\]That's your first step ya? :D

zepdrix (zepdrix):

you're* that's gonna bug me, i had to lol

OpenStudy (anonymous):

yeah i even crossed it out and all i guess it just slipped my mind when i was rewriting it on the other side of the equal sign

OpenStudy (anonymous):

anyway so its \[y'= \frac{ 4x^3 + 36x + 2y }{ 2(x+y) }\]

zepdrix (zepdrix):

Woops, -2y on top I think ya?

zepdrix (zepdrix):

Anyway, let's just get rid of all the 2's I guess,\[\large\rm y'=\frac{2x^3+18x-y}{x+y}\]That's the only simplification that really cleans it up nicely.

OpenStudy (anonymous):

:( yes , okay so now what?

zepdrix (zepdrix):

This derivative function is undefined when the denominator is zero. (This is also when we're getting vertical tangents.)

zepdrix (zepdrix):

So vertical tangent when the denominator is zero, \(\large\rm x+y=0\)

zepdrix (zepdrix):

Overheat again? :) LOL

OpenStudy (jhannybean):

 where its undefined? :P

OpenStudy (anonymous):

yep :( i need a new computer

zepdrix (zepdrix):

Hmm ya that's a weird answer :o I do something wrong?

OpenStudy (anonymous):

anyway one of the answer choices is x = -y so i think that's the answer right?

zepdrix (zepdrix):

yay team \c:/ it just doesn't make a whole lot of sense with the graph of the function :D I guess I just need to think about it a sec lol

OpenStudy (jhannybean):

I think that's about right, it's either \(\sf y=-x\) or \(\sf x=-y\).

OpenStudy (anonymous):

thank you once more :D